OpenStudy (anonymous):

a 80 meter cable is cut in 2pcs the length of pcs have ratio of 11to 5 . how long is shorter pc. i have 15x divided by 80 =5.3 11times 5.3 +5times5.3 but it doesnt add up please explain step by step how to solve its driving me crazy

OpenStudy (anonymous):

dude, you should 've 16x and not 15x.... 11+5=16

OpenStudy (anonymous):

11x and 5x forming 80 5x =( 80/16) * 5 = 25 shorter is 25, longer is 55